OpenStudy (anonymous):

if a dress is on sale for 25% off, what percent of the cost does the customer pay ? show work pls the dress is 45 dollars

jhonyy9 (jhonyy9):

45 ----- 100% x ----- 25% ------------- x= 45*25/100 = 45/4 = 11,25 45-11,25 = 33,75

jhonyy9 (jhonyy9):

ok ? the customer will pay 33,75 dollars

OpenStudy (anonymous):

thank uu

jhonyy9 (jhonyy9):

was my pleasure